# Angular momentum of vacuum bubbles in a first-order phase transition

### v1

@@ 3 Angular momentum from cosmological perturbations
In our approach, we would like to estimate the angular momentum of a volume enclosing a dark fluid, induced by cosmological perturbations. We assume that these volumes are spherical to simplify our calculations. To set up the starting expression for the angular momentum, we first provide a discussion of some well-known subtleties regarding its definition in GR (cf. [28, 29, 30, 31, 32]).
@@ 3.1 Angular momentum in general relativity
Furthermore, one can show that the total mass and angular momentum reduce into surface integrals. This follows from the linearized field equations, expanded about Minkowski spacetime, which can be written as [29]
@@
[…]
Note that even in the linearized case, the integrands of the flux integrals are gauge-dependent, because we have the freedom to choose a coordinate system where the derivatives of the metric vanish. However, the surface integrals are defined at the boundary where the spacetime is Minkowski, and hence the quantities $P^{\mu}$ and $J^{\mu\nu}$ behave as tensors, and this presents no ambiguity regarding the values of $M$ and $J^{c}$. In a generic spacetime, however, the energy momentum tensor $T^{\mu\nu}$ is covariantly conserved, i.e.
@@
where $\nabla_{\mu}$ is the covariant derivative associated with the Christoffel connection $\Gamma^{\nu}_{\alpha\beta}$. This does not automatically imply the existence of a global conservation law. Intuitively, this can be traced back to the fact that the gravitational field itself, not just the matter fields, generates energy. Heuristically, the contribution of the gravitational field to the total energy momentum is expected to involve a product of metric derivatives, by analogy with the Newtonian case where the energy density is the square of the gradient of the gravitational potential. Then one can construct a quantity $\tau_{\mu\nu}$, which represents the stress energy of the gravitational field, such that, when combined with the matter stress energy tensor $T^{\mu\nu}$, yields a global conservation law. As an example, Landau and Lifshitz [28] constructed (see also [33] for a unified treatment in terms of superpotentials [34, 35, 36, 37, 38]; not to be confused with the superpotential in supersymmetry)
@@
where $g_{3}$ is the determinant of the induced metric on $\Sigma$. We note that the angular momentum in Eq. (3.15) was adopted in [6] to calculate the spin of PBHs in a flat FRW background. In the case of a perturbed FRW background however, there is no Killing vector associated with rotational symmetry to begin with. Nevertheless, globally conserved quantities can still be defined in the case of perturbed spacetimes.
@@ 3.2 Angular momentum in a perturbed FRW spacetime
The discussion in [39] will be useful for our task in determining the angular momentum of fluids subjected to cosmological perturbations, where they constructed conserved charges based on the seminal work [37]. Reference [39] formulated a procedure to obtain superpotentials and conserved quantities, followed by an application of the formalism in the case of perturbed FRW backgrounds. We refer the interested reader to [39] for the explicit expressions for the superpotentials and conserved quantities in a general curved spacetime, and instead we only quote the relevant results. Consider a flat FRW background endowed with a metric $\bar{g}_{\mu\nu}=a^{2}(\eta)\eta_{\mu\nu}$, and $\xi$ are chosen to be the 15 conformal Killing vectors of Minkowski spacetime—4 spacetime translations, 3 from spatial rotations, 3 from Lorentz boosts, 1 from dilation and 4 from special conformal transformations. Taking the time component of the conserved quantity $\hat{I}^{\mu}$, with an associated superpotential $\hat{I}^{\mu\nu}$, we have
@@
Note that Eqs. (3.17)-(3.20) are exact expressions, even for “large” perturbations of the background FRW metric $\bar{g}$. Note that the Latin indices are raised and lowered using the Kronecker delta. Typically, we are interested in the volume integral of $\hat{I}^{0}$ over a spherical volume at a given conformal time $\eta$. Ref. [39] demonstrated that the conserved charges are gauge dependent, and these quantities take on simpler forms in the uniform Hubble gauge [40, 41, 21, 42], where the change in the trace of the extrinsic curvature of surfaces of constant conformal time, under metric perturbations, is zero. In this gauge, the perturbed FRW metric in comoving coordinates can be written as [40, 21]
@@
We emphasize here that the angular momentum obtained in Eq. (3.26), through the formalism of [39], coincides with the starting expression in [6, 7]. The latter obtained Eq. (3.26) by assuming that the hypersurface integral version of the Komar formula for angular momentum, Eq. (3.15), still holds even in the perturbed FRW case by simply replacing $T_{\mu\nu}$ with the perturbed energy momentum tensor.
@@
[…]
It was pointed out by [43, 5] that one must consider the second-order contribution to angular momentum in the case of spherical volumes. The determinant of the spatial part of the metric, to first order in perturbations, is given by
@@
where $\delta$ is the fluid density contrast, and $v_{j}$ is a dimensionless velocity term, which arises from a potential, i.e.
@@
Note that the integrand consists of products of two first-order quantities. Aside from the product of the density contrast and fluid velocity, which is present in [43] and [5], we also have the following contributions: the products of the fluid velocity and gravitational potential, and the fluid velocity and $\nabla^{2}\gamma$.
@@
[…]
@@
$R(\eta)=x_{0}a(\eta)\,,$ $d\Pi_{k}\equiv\frac{d^{3}\vec{k}}{(2\pi)^{3}}\,,\quad\mathcal{F}(z)\equiv\frac{3z\cos z-(3-z^{2})\sin z}{z^{5}}\,.$ (3.36)
The quantity $R$ is the physical radius of the false vacuum bubble with comoving size $x_{0}$, at conformal time $\eta$. Following [43], we subtract the contribution from the translational motion of the sphere’s center of mass (CM). We find that the angular momentum about the CM is
[…]
@@
[…]
×$\times\int_{-1}^{1}dx~{}(1-x^{2})~{}W^{2}(z,z^{\prime};\tilde{x}_{0})\Delta_{\mathcal{R}}^{2}(k)\Delta_{\mathcal{R}}^{2}(k^{\prime})\,,$ (3.43)
@@
We have also introduced the rescaled form of the curvature power spectrum
@@
where $A_{\rm s}$ is the amplitude of superhorizon curvature perturbations, $k_{\rm s}$ is the tilt scale, and $n_{\rm s}$ is the spectral index of scalar perturbations. The measured values are given by [44]:
@@
In Eq. (3.43) it is understood that, e.g. $\mathcal{F}_{k+k^{\prime}}\equiv\mathcal{F}(|\vec{k}+\vec{k}^{\prime}|x_{0})$, $\mathcal{F}_{k}\equiv\mathcal{F}(kx_{0})$ and $\mathcal{G}_{k}\equiv\mathcal{G}(kx_{0})$. We have introduced the parameter,
@@ 3.3 Spin of primordial black holes
Our calculation of the angular momentum and spin of false vacuum remnants is motivated by the desire to determine the spin of PBHs formed by the direct collapse of FV bubbles or by the collapse of Fermi balls. On the other hand, the spin of PBHs from cosmological perturbations has been discussed extensively in the literature, mainly in Refs. [6, 5, 7], in the usual context of PBH formation through the collapse of overdensities. These sizable density perturbations in the cosmological fluid are seeded by rare peaks in the curvature power spectrum. Such enhancements in the power spectrum may be realized in certain inflationary models [45, 46, 47]. However, there seems to be an apparent disagreement between the two main approaches to this mechanism. Reference [6] argued that it is appropriate to consider ellipsoidal volumes in describing overdensities. This follows from an earlier work by [16] in which the overdensities are expanded to quadratic order about the peak and the overdensity within the volume is required to be at least a certain fraction of the peak value, which is one of the criteria to ensure the collapse into PBHs; then, in general, the boundary of this volume is triaxial. The relevant phase space of triaxial configurations and velocity shear has 16 dimensions, where the phase space variables are: overdensity; first and second derivatives of the overdensity; and the first derivative of the velocity field. The joint probability distribution over this configuration space is essentially Gaussian [48], where the covariance matrix, which contains the two point correlators of the phase space variables, is ultimately determined by the fluid transfer functions and curvature perturbations. In choosing the gauge, they considered constant mean curvature slicing [49, 17, 50], which is equivalent to the uniform Hubble gauge, and worked in the subhorizon limit for the fluid perturbations. By considering ellipsoidal volumes, [6] showed that, starting from the Komar expression, angular momentum is a quantity that is first order in perturbations.
@@
In contrast, the authors of Ref. [5] considered spherical volumes in their PBH spin calculation, which they have justified by referring to [48]. Firstly, they assumed that curvature perturbations $\zeta$ have rare high peaks in position space. Once these perturbations enter the horizon, they collapse into PBHs. In position space, one can construct contours of constant curvature, as shown in Figure 1 of Ref. [5]. The ellipticity of those contours is proportional to $\zeta_{\rm rms}/\zeta_{0}$, where $\zeta_{0}$ is the amplitude of the peak. Sending $\zeta_{\rm rms}\rightarrow 0$ brings the ellipticity to zero. The angular momentum is calculated using Eq. (3.10), at a time when the PBH has already formed. It was argued, in a manner similar to Eq. (3.28), that the first-order contribution to the angular momentum is zero for a spherical volume, and thus the leading nonvanishing contribution is at second order in perturbations. The size of the integration volume is chosen carefully: the 3-volume around the PBH must be sufficiently large such that the spacetime geometry generated by the PBH is close to being Minkowski, but small enough that it does not see the curvature due to Hubble. These criteria can be easily realized at late times, but still during the radiation dominated era of the Universe, where the Hubble horizon has grown to a larger size. They then calculated the flux of angular momentum from acoustic waves through this 3-volume, and identified it with the torque on the PBH. Because they were interested in the initial spin of PBHs, they extrapolated their result to early times. It is worth noting that they worked in the Newtonian gauge, used the subhorizon limit for the evolution of perturbations, and considered an enhanced power spectrum with a narrow Gaussian peak.
@@ 3.4 Angular momentum due to uniform pressure
Here we have taken $p$ to be the uniform pressure on the surface. The above expression can be shown to be a total derivative, i.e.
@@ 4 Evolution of cosmological perturbations
The angular momentum of a spherical volume containing several fluid components can be computed once the perturbations on each fluid component and the gravitational potentials $\Phi$ and $\gamma$ are known. We identify the periods in cosmological history that are crucial for us: the conformal time $\eta=\eta_{0}\simeq 0$ at which all perturbations are superhorizon, the critical point $\eta=\eta_{\rm c}$, and the onset of percolation $\eta=\eta_{*}$. In what follows, quantities evaluated at the percolation time are indicated with a subscript *. Here we aim to calculate only the perturbations, and eventually the angular momentum, during $[0,\eta_{\rm c}]$ and $[\eta_{\rm c},\eta_{*}]$. The subsequent evolution, which will result in either the direct collapse to a PBH or the formation of a FB before collapsing into a PBH, is described by different dynamics beyond the scope of our study.
Following Ref. [21], and working in the UHG, we can track the evolutions of $\delta$, $\hat{\psi}$, and $\Phi$. For a fluid component $f$, the equation of state and sound speeds, defined as
$w_{f}\equiv\frac{p_{f}}{\rho_{f}}\,,\quad c_{s,f}^{2}\equiv\left(\frac{\partial\rho_{f}}{\partial p_{f}}\right)_{s}=\frac{d\rho_{f}/dT}{dp_{f}/dT}\,,$ (4.1)
@@
where we used the Friedmann equation in the last equation. Meanwhile, the metric perturbation $\gamma$ in Fourier space can be obtained from

### v2

@@ 3 Angular momentum from cosmological perturbations
In our approach, we would like to estimate the angular momentum of a volume enclosing a dark fluid, induced by cosmological perturbations. We assume that these volumes are spherical to simplify our calculations. To set up the starting expression for the angular momentum, we first provide a discussion of some well-known subtleties regarding its definition in GR (cf. [landau1975classical, thorne2000gravitation, Wald:1984rg, Poisson:2009pwt, Jaramillo:2010ay]).
@@ 3.1 Angular momentum in general relativity
Furthermore, one can show that the total mass and angular momentum reduce into surface integrals. This follows from the linearized field equations, expanded about Minkowski spacetime, which can be written as [thorne2000gravitation]
@@
$J^{\mu\nu}$ $=$ $\frac{1}{16\pi G_{\rm N}}\oint_{\partial\mathcal{V}}\left(x^{\mu}\partial_{\beta}H^{\nu\beta 0j}-x^{\nu}\partial_{\beta}H^{\mu\beta 0j}+H^{\mu j0\nu}-H^{\nu j0\mu}\right)n_{j}dS\,.$ (3.6)
Note that even in the linearized case, the integrands of the flux integrals are gauge-dependent, because we have the freedom to choose a coordinate system where the derivatives of the metric vanish. However, the surface integrals are defined at the boundary where the spacetime is Minkowski, and hence the quantities $P^{\mu}$ and $J^{\mu\nu}$ behave as tensors, and this presents no ambiguity regarding the values of $M$ and $J^{c}$. In a generic spacetime, however, the energy momentum tensor $T^{\mu\nu}$ is covariantly conserved, i.e.,
@@
where $\nabla_{\mu}$ is the covariant derivative associated with the Christoffel connection $\Gamma^{\nu}_{\alpha\beta}$. This does not automatically imply the existence of a global conservation law. Intuitively, this can be traced back to the fact that the gravitational field itself, not just the matter fields, generates energy. Heuristically, the contribution of the gravitational field to the total energy momentum is expected to involve a product of metric derivatives, by analogy with the Newtonian case where the energy density is the square of the gradient of the gravitational potential. Then one can construct a quantity $\tau_{\mu\nu}$, which represents the stress energy of the gravitational field, such that, when combined with the matter stress energy tensor $T^{\mu\nu}$, yields a global conservation law. As an example, Landau and Lifshitz [landau1975classical] constructed (see also [trautman1962conservation] for a unified treatment in terms of superpotentials [ph1939uber, Goldberg:1953zz, Komar:1958wp, Katz:1996nr, Chang:1998wj]; not to be confused with the superpotential in supersymmetry)
@@
where $g_{3}$ is the determinant of the induced metric on $\Sigma$. We note that the angular momentum in Eq. (3.16) was adopted in [DeLuca:2019buf] to calculate the spin of PBHs in a flat FRW background. In the case of a perturbed FRW background however, there is no Killing vector associated with rotational symmetry to begin with. Nevertheless, globally conserved quantities can still be defined in the case of perturbed spacetimes.
@@ 3.2 Angular momentum in a perturbed FRW spacetime
The discussion in Ref. [Petrov:1999tbb] will be useful for our task in determining the angular momentum of fluids subjected to cosmological perturbations, where they constructed conserved charges based on the seminal work in Ref. [Katz:1996nr]. Reference [Petrov:1999tbb] formulated a procedure to obtain superpotentials and conserved quantities, followed by an application of the formalism in the case of perturbed FRW backgrounds. We refer the interested reader to Ref. [Petrov:1999tbb] for the explicit expressions for the superpotentials and conserved quantities in a general curved spacetime, and instead we only quote the relevant results. Consider a flat FRW background endowed with a metric $\bar{g}_{\mu\nu}=a^{2}(\eta)\eta_{\mu\nu}$, and $\xi$ are chosen to be the 15 conformal Killing vectors of Minkowski spacetime—4 spacetime translations, 3 from spatial rotations, 3 from Lorentz boosts, 1 from dilation and 4 from special conformal transformations. Taking the time component of the conserved quantity $\hat{I}^{\mu}$, with an associated superpotential $\hat{I}^{\mu\nu}$, we have
@@
Note that Eqs. (3.17)-(3.20) are exact expressions, even for “large” perturbations of the background FRW metric $\bar{g}$. Note that the Latin indices are raised and lowered using the Kronecker delta. Typically, we are interested in the volume integral of $\hat{I}^{0}$ over a spherical volume at a given conformal time $\eta$. Ref. [Petrov:1999tbb] demonstrated that the conserved charges are gauge dependent, and these quantities take on simpler forms in the uniform Hubble gauge [Bardeen:1980kt, kodama1984cosmological, Schmid:1998mx, hwang1999relativistic], where the change in the trace of the extrinsic curvature of surfaces of constant conformal time, under metric perturbations, is zero. In this gauge, the perturbed FRW metric in comoving coordinates can be written as [Bardeen:1980kt, Schmid:1998mx]
@@
We emphasize here that the angular momentum obtained in Eq. (3.26), through the formalism of [Petrov:1999tbb], coincides with the starting expression in [DeLuca:2019buf, Harada:2020pzb]. The latter obtained Eq. (3.26) by assuming that the hypersurface integral version of the Komar formula for angular momentum, Eq. (3.16), still holds even in the perturbed FRW case by simply replacing $T_{\mu\nu}$ with the perturbed energy momentum tensor.
@@
$\int_{\mathcal{V}}\epsilon^{kal}x_{l}\partial_{k}\psi~d^{3}\vec{x}=\oint_{\partial\mathcal{V}}\epsilon^{kal}x_{0}n_{l}n_{k}\psi~dS=0\,.$ (3.28)
It was pointed out by [Peebles:1969jm, Mirbabayi:2019uph] that one must consider the second-order contribution to angular momentum in the case of spherical volumes. The determinant of the spatial part of the metric, to first order in perturbations, is given by
@@
where $\delta$ is the fluid density contrast, and $v_{j}$ is a dimensionless velocity term, which arises from a potential, i.e.,
@@
Note that the integrand consists of products of two first-order quantities. Aside from the product of the density contrast and fluid velocity, which is present in [Peebles:1969jm] and [Mirbabayi:2019uph], we also have the following contributions: the products of the fluid velocity and gravitational potential, and the fluid velocity and $\nabla^{2}\gamma$.
@@
$\vec{J}$ $=$ $-a~4\pi\bar{\rho}R^{4}x_{0}\int d\Pi_{k}d\Pi_{k^{\prime}}(\vec{k}\times\vec{k}^{\prime})~\delta_{\rm eff}(\vec{k})~\frac{\hat{\psi}(\vec{k}^{\prime})}{k^{\prime}}~\mathcal{F}(|\vec{k}+\vec{k}^{\prime}|x_{0})\,$, (3.34)
@@
$R(\eta)$ $=x_{0}a(\eta)\,,\quad d\Pi_{k}\equiv\frac{d^{3}\vec{k}}{(2\pi)^{3}}\,,\quad\mathcal{F}(z)\equiv\frac{3z\cos z-(3-z^{2})\sin z}{z^{5}}\,.$ (3.36)
The quantity $R$ is the physical radius of the false vacuum bubble with comoving size $x_{0}$, at conformal time $\eta$. Following [Peebles:1969jm], we subtract the contribution from the translational motion of the sphere’s center of mass (CM). We find that the angular momentum about the CM is
$\vec{J}_{\rm CM}(\eta)$ $=$ $4\pi\bar{\rho}R^{5}\int d\Pi_{k}d\Pi_{k^{\prime}}~(\vec{k}\times\vec{k}^{\prime})~\delta_{\rm eff}(\vec{k},\eta)$ (3.37)
@@
$\mathcal{C}(\tilde{x}_{0},x_{0})$ $\equiv A_{\rm s}^{-2}\int_{0}^{\infty}dz~z\int_{0}^{\infty}dz^{\prime}~z^{\prime}$
$\times\int_{-1}^{1}dx~(1-x^{2})~W^{2}(z,z^{\prime};\tilde{x}_{0})\Delta_{\mathcal{R}}^{2}(k)\Delta_{\mathcal{R}}^{2}(k^{\prime})\,$, (3.43)
@@
We have also introduced the rescaled form of the curvature power spectrum,
@@
where $A_{\rm s}$ is the amplitude of superhorizon curvature perturbations, $k_{\rm s}$ is the tilt scale, and $n_{\rm s}$ is the spectral index of scalar perturbations. The measured values are given by [Planck:2018vyg]:
@@
[…]
@@ 3.3 Spin of primordial black holes
Our calculation of the angular momentum and spin of false vacuum remnants is motivated by the desire to determine the spin of PBHs formed by the direct collapse of FV bubbles or by the collapse of Fermi balls. On the other hand, the spin of PBHs from cosmological perturbations has been discussed extensively in the literature, mainly in Refs. [DeLuca:2019buf, Mirbabayi:2019uph, Harada:2020pzb], in the usual context of PBH formation through the collapse of overdensities. These sizable density perturbations in the cosmological fluid are seeded by rare peaks in the curvature power spectrum. Such enhancements in the power spectrum may be realized in certain inflationary models [Bartolo:2018rku, Clesse:2018ogk, Vaskonen:2020lbd]. However, there seems to be an apparent disagreement between the two main approaches to this mechanism. Reference [DeLuca:2019buf] argued that it is appropriate to consider ellipsoidal volumes in describing overdensities. This follows from an earlier work by [heavens1988tidal] in which the overdensities are expanded to quadratic order about the peak and the overdensity within the volume is required to be at least a certain fraction of the peak value, which is one of the criteria to ensure the collapse into PBHs; then, in general, the boundary of this volume is triaxial. The relevant phase space of triaxial configurations and velocity shear has 16 dimensions, where the phase space variables are: overdensity; first and second derivatives of the overdensity; and the first derivative of the velocity field. The joint probability distribution over this configuration space is essentially Gaussian [Bardeen:1985tr], where the covariance matrix, which contains the two point correlators of the phase space variables, is ultimately determined by the fluid transfer functions and curvature perturbations. In choosing the gauge, they considered constant mean curvature slicing [Shibata:1999zs, Harada:2013epa, Harada:2015yda], which is equivalent to the uniform Hubble gauge, and worked in the subhorizon limit for the fluid perturbations. By considering ellipsoidal volumes, [DeLuca:2019buf] showed that, starting from the Komar expression, angular momentum is a quantity that is first order in perturbations.
@@
In contrast, the authors of Ref. [Mirbabayi:2019uph] considered spherical volumes in their PBH spin calculation, which they have justified by referring to [Bardeen:1985tr]. Firstly, they assumed that curvature perturbations $\zeta$ have rare high peaks in position space. Once these perturbations enter the horizon, they collapse into PBHs. In position space, one can construct contours of constant curvature, as shown in Figure 1 of Ref. [Mirbabayi:2019uph]. The ellipticity of those contours is proportional to $\zeta_{\rm rms}/\zeta_{0}$, where $\zeta_{0}$ is the amplitude of the peak. Sending $\zeta_{\rm rms}\rightarrow 0$ brings the ellipticity to zero. The angular momentum is calculated using Eq. (3.10), at a time when the PBH has already formed. It was argued, in a manner similar to Eq. (3.28), that the first-order contribution to the angular momentum is zero for a spherical volume, and thus the leading nonvanishing contribution is at second order in perturbations. The size of the integration volume is chosen carefully: the 3-volume around the PBH must be sufficiently large such that the spacetime geometry generated by the PBH is close to being Minkowski, but small enough that it does not see the curvature due to Hubble. These criteria can be easily realized at late times, but still during the radiation dominated era of the Universe, where the Hubble horizon has grown to a larger size. They then calculated the flux of angular momentum from acoustic waves through this 3-volume, and identified it with the torque on the PBH. Because they were interested in the initial spin of PBHs, they extrapolated their result to early times. It is worth noting that they worked in the Newtonian gauge, used the subhorizon limit for the evolution of perturbations, and considered an enhanced power spectrum with a narrow Gaussian peak.
@@ 3.4 Angular momentum due to uniform pressure
Here we have taken $p$ to be the uniform pressure on the surface. The above expression can be shown to be a total derivative, i.e.,
@@ 4 Evolution of cosmological perturbations
The angular momentum of a spherical volume containing several fluid components can be computed once the perturbations on each fluid component and the gravitational potentials $\Phi$ and $\gamma$ are known. We identify the periods in cosmological history that are crucial for us: the conformal time $\eta=\eta_{0}\simeq 0$ at which all perturbations are superhorizon, the critical point $\eta=\eta_{\rm c}$, and the onset of percolation $\eta=\eta_{*}$. In what follows, quantities evaluated at the percolation time are indicated with a subscript *. Here we aim to calculate only the perturbations, and eventually the angular momentum, during $0,\eta_{\rm c}]$ and $\eta_{\rm c},\eta_{*}]$. The subsequent evolution, which will result in either the direct collapse to a PBH or the formation of a FB before collapsing into a PBH, is described by different dynamics beyond the scope of our study.
Following Ref. [Schmid:1998mx], and working in the UHG, we can track the evolutions of $\delta$, $\hat{\psi}$, and $\Phi$. For a fluid component $f$, the equation of state and sound speeds, defined as
$w_{f}\equiv\frac{p_{f}}{\rho_{f}}\,,\quad c_{s,f}^{2}\equiv\left(\frac{\partial p_{f}}{\partial\rho_{f}}\right)_{s}=\frac{dp_{f}/dT}{d\rho_{f}/dT}\,$, (4.1)
@@
where we used the Friedmann equation in the last equation. The metric perturbation $\gamma$ in Fourier space can be obtained from